# Theory of dynamical superradiance in organic materials

### v1

@@ Appendix A Influence of $\theta$ on dynamical SR
In this appendix, we investigate the influence of the angle $\theta$ on dynamical SR in the presence of vibrational dressing. All results are calculated with mean-field theory. Figure 7 shows the average number of photons and the electronic coherence as a function of time for three different values of $\theta$ [Figs. 7(a) and 7(d) are the same as Figs. 4(a) and 4(b)]. Clearly, an increase in $\theta$ leads to an increase in initial electronic coherence $|\langle\sigma^{+}\rangle|=\sin(\theta)/2$, resulting in an increased photon emission rate. However, irrespective of $\theta$, for strong vibrational coupling $S\gtrsim 0.4$, the coherence tends to decrease at early times, which indicates that SR is suppressed.
@@
Figures 8(a) and 8(b) show how the angle $\theta$ changes the risetime behavior in the case of vibrational dressing for $\Delta=0$ and $\Delta=-0.3\,\mathrm{eV}$, respectively. We set $S=0.2$, which ensures that the system still exhibits an exponential rise behavior at early times. We see that for small values of $\theta$, the risetime (defined by fitting $A\mathrm{e}^{t/\tau}$ to the exponential regions) is approximately constant. However, increasing $\theta$ results in an enhanced initial photon emission, before the exponential regime is reached. If $\theta$ becomes too large, the exponential regime disappears.

### v2

@@ Appendix B Influence of $\theta$ on dynamical SR
In this appendix, we investigate the influence of the angle $\theta$ on dynamical SR in the presence of vibrational dressing. All results are calculated with mean-field theory. Figure 8 shows the average number of photons and the electronic coherence as a function of time for three different values of $\theta$ [Figs. 8(a) and 8(d) are the same as Figs. 4(a) and 4(b)]. Clearly, an increase in $\theta$ leads to an increase in initial electronic coherence $|\langle\hat{\sigma}^{+}\rangle|=\sin(\theta)/2$, resulting in an increased photon emission rate. However, irrespective of $\theta$, for strong vibrational coupling $S\gtrsim 0.4$, the coherence tends to decrease at early times, which indicates that SR is suppressed.
@@
Figures 9(a) and 9(b) show how the angle $\theta$ changes the risetime behavior in the case of vibrational dressing for $\Delta=0$ and $\Delta=-0.3\,\mathrm{eV}$, respectively. We set $S=0.2$, which ensures that the system still exhibits an exponential rise behavior at early times. We see that for small values of $\theta$, the risetime (defined by fitting $A\mathrm{e}^{t/\tau}$ to the exponential regions) is approximately constant. However, increasing $\theta$ results in an enhanced initial photon emission, before the exponential regime is reached. If $\theta$ becomes too large, the exponential regime disappears.
@@
Appendix C Influence of $g\sqrt{N}$ on dynamical SR
We study here in more detail the effect of the light-matter coupling strength $g\sqrt{N}$ on the threshold of dynamical SR reported in Sec. III. To this end, we calculate the same dynamics as in Figs. 4(a) and 4(b) in the main text for three different values of $g\sqrt{N}$, while all other parameters are kept the same. The results are shown in Fig. 10, where we note that Figs. 10(c) and 10(d), corresponding to $g\sqrt{N}=0.2\,\mathrm{eV}$, are the same as Figs. 4(a) and 4(b). Recall that for $g\sqrt{N}=0.2\,\mathrm{eV}$, we identified a threshold for dynamical SR for $S\approx 0.3$, since beyond that value of $S$, the electronic coherence does not monotonically increase anymore. From Figs. 10(b) and 10(f), we see that the so-defined SR threshold is at roughly $S\approx 0.1$ for $g\sqrt{N}=0.05\,\mathrm{eV}$, whereas for $g\sqrt{N}=0.4\,\mathrm{eV}$, even for $S=0.8$ the electronic coherence rises monotonically. We therefore conclude that the SR threshold emerges from a competition between light-matter coupling and vibronic coupling: The stronger the light-matter coupling, the larger the maximum value of $S$ for which dynamical SR is still possible. This makes sense, since $g\sqrt{N}$ corresponds to the Rabi-oscillation frequency of the cavity-emitter system, i.e. larger values of $g\sqrt{N}$ result in a faster initial emission of photons, as seen in Figs. 10(a), 10(c) and 10(e).
Figure 10: Mean-field solution of the average number of photons $\langle\hat{n}\rangle/N=|\langle\hat{a}\rangle|^{2}/N$ and the electronic coherence $|\langle\hat{\sigma}^{+}\rangle|$ in the HTC model for different values of $S$ and $g\sqrt{N}$. (a)-(b): $g\sqrt{N}=0.05\,\mathrm{eV}$, (c)-(d): $g\sqrt{N}=0.2\,\mathrm{eV}$ [same as Fig. 4(a)-(b)], (e)-(f): $g\sqrt{N}=0.4\,\mathrm{eV}$. All other parameters are the same as in Fig. 4(a)-(c).